# Large Language Models Miss the Multi-Agent Mark

### v2

@@ 4 Coordination and Communication Issues in MAS LLMs
Coordination and the lack of asynchronous MAS LLMs. As described in Section 1,margin: 1 margin: the benefits of
asynchronicity
asynchronicity is a key component of multi-agent concurrent systems: in its basic form, concurrent systems encompass multiple, diverse tasks executed simultaneously, without assuming when they start or end. Examples of notorious problems that can only be modelled with asynchronicity are those that necessitate concurrent algorithms to be solved (10.5555/2029110, ; diaz1981formalization, ): these include classic problems such as the “dining philosophers”, handshaking protocols, etc. Asynchronicity also arises in many practical scenarios, such as email and chat exchanges and database management access. Many real-world scenarios cannot be modelled without asynchronicity and require simplification (e.g., by assuming agents act sequentially through an orchestrator).
Notably, while most closed-source LLM providers offer asynchronous APIs,margin: 1 margin: LLMs are
asynchronous…
agents employing LLMs tend to be predominantly used in synchronous or parallel fashion (li2024agentsneed, ).
In this sense, we surveyed the MAS LLMs literaturemargin: 1 margin: their interactions
are not!
published between $2022$ and $2025$, to understand how many works directly employ asynchronicity or enable the deployment of asynchronous agents. We identified few works ($22$) that explicitly model or discuss asynchronous agent interactions. [^5] As a reference, this survey counts more than $1400$ MAS LLMs articles: https://github.com/AGI-Edgerunners/LLM-Agents-Papers Furthermore, in those few cases, asynchronous interactions are implemented through conversations and by employing frameworks and languages that are not natively asynchronous, which adds unnecessary complexity and reduces interoperability. More information is provided in Appendix B.
@@
Communications methods in MAS LLMs. Traditionally, there are three levels at which (Multi-Agent) communication is analysed (austin1962how, ). An utterance that conveys some meaning may have to the hearer no intended effect (an illocutionary act, e.g., "the sky is blue"),margin: 1 margin: agent
communication
principles
some meaning intended to warn (a perlocutionary act, "a train is passing"), or some meaning that acts as a request (a performative, e.g., "please open the window"). These distinctions, alongside the so-called rules of conversation (implicature, the Gricean maxims (Grice1975-GRILAC-6, ), etc.), constitute the foundation of agents’ communication in MAS (berna2004communication, ; jang2005efficient, ; saunders1996evolution, ).
In MAS,margin: 1 margin: performative
communication
illocutionary conversations are usually handled using descriptive and structured languages, such as JSON and RDF. Their purpose is to exchange information, not to perform or request actions. On the other hand, KQML (kqml1994, ), FIPA’s ACL, [^8] http://www.fipa.org/specs/fipa00061/XC00061D.html and rational programming and Agent-Oriented Programming (shoham1993agent, ) are examples of consolidated performative standards that describe agents’ beliefs, commitments, and actions. For instance, for an agent that exchanges a file with another and asks to summarise it, the MAS literature proposes using a structured language for the exchange and a performative query for the summarisation.
When it comes to LLMs,margin: 1 margin: natural language
is ambiguous
humans interact with them in free-text form; by extension, most MAS LLMs systems adopt natural language as the primary communication medium. While natural language captures the nuances of the principles mentioned above, it comes at the cost of complexity and ambiguity: an utterance such as "a car is coming" may range from being illocutionary (conveying information) to performative (conveying an implicit warning or an order). A few recent works in MAS LLMs propose to handle handshakes and errors with natural language communications; any other communication that routines and protocols can implement should otherwise favour structured languages (marro2024scalablecommunicationprotocolnetworks, ).
@@ 5 MAS LLMs do not Quantify the Emergence of Complex Behaviours
In many disciplines,margin: 1 margin: emergence such as system theory, economics, MAS, etc., an emergent behaviour is observed when a complex entity has properties or behaviours that its parts do not have on their own, and emerge only when they interact in a wider whole (altmann2024emergencemultiagentsystemssafety, ; fromm2005types, ). In the context of LLMs, emergence describes the increasing capabilities of a model at varying model size (brown2020languagemodelsfewshotlearners, ; kaplan2020scalinglawsneurallanguage, ). With MAS LLMs, behaviours that cannot be predicted via a static analysis of agents and their environment are also considered emergent Li_2023 ; park2023generative .
Emergent behaviours in MAS and MAS LLMsmargin: 1 margin: observational
emergence
are often associated with open-ended environments, i.e., those MAS settings where LLMs interact and evolve freely. In an influential work park2023generative , Park et al. use LLMs to simulate a sandboxed society, with a focus on how single instructions influence the population and information spreads across the environment. While the authors show that simply nudging one agent causes other agents to engage in complex behaviours, the concept of emergence itself is never addressed formally.
@@
We thus surveyedmargin: 1 margin: emergence in
MAS LLMs
papers published in the MAS LLMs and machine learning community between $2023$ and $2025$ that mention emergent behaviours to understand what methods and metrics employ to identify and quantify them. Out of more than $60$ papers analysed, only a few define clear metrics to measure emergent behaviours, while the majority qualitatively evaluate such behaviours and report the most notable. More details about the analysis and the list of papers are reported in Appendix C.
@@
In conclusion,margin: 1 margin: quantifying
emergence
in this body of research, we observe (i) no systematic definition of what constitutes an emergent behaviour, without reference to the system being analysed, therefore (ii) a lack of proper benchmarks to quantify them. In contrast, traditional MAS research generally insists on rigorously defined environments, quantifiable objectives, and formal verification methods. Emergence alone, especially when derived from loosely constrained prompts and undefined interaction dynamics, does not suffice to claim long-horizon MAS-level coordination or planning capabilities (turner2022formalizingproblemeffectregularization, ).

### v3

@@ 4 Coordination and Communication Issues in MAS LLMs
Coordination and the lack of asynchronous MAS LLMs. As described in Section 1, asynchronicity is a key component of multi-agent concurrent systems: in its basic form, concurrent systems encompass multiple, diverse tasks executed simultaneously, without assuming when they start or end. Examples of notorious problems that can only be modelled with asynchronicity are those that necessitate concurrent algorithms to be solved (10.5555/2029110, ; diaz1981formalization, ): these include classic problems such as the “dining philosophers”, handshaking protocols, etc. Asynchronicity also arises in many practical scenarios, such as email and chat exchanges and database management access. Many real-world scenarios cannot be modelled without asynchronicity and require simplification (e.g., by assuming agents act sequentially through an orchestrator).
Notably, while most closed-source LLM providers offer asynchronous APIs,.
In this sense, we surveyed the MAS LLMs literature published between $2022$ and $2025$, to understand how many works directly employ asynchronicity or enable the deployment of asynchronous agents. We identified few works ($22$) that explicitly model or discuss asynchronous agent interactions. [^5] As a reference, this survey counts more than $1400$ MAS LLMs articles: https://github.com/AGI-Edgerunners/LLM-Agents-Papers Furthermore, in those few cases, asynchronous interactions are implemented through conversations and by employing frameworks and languages that are not natively asynchronous, which adds unnecessary complexity and reduces interoperability. More information is provided in Appendix B.
@@
Communications methods in MAS LLMs. Traditionally, there are three levels at which (Multi-Agent) communication is analysed (austin1962how, ). An utterance that conveys some meaning may have to the hearer no intended effect (an illocutionary act, e.g., "the sky is blue"), some meaning intended to warn (a perlocutionary act, "a train is passing"), or some meaning that acts as a request (a performative, e.g., "please open the window"). These distinctions, alongside the so-called rules of conversation (implicature, the Gricean maxims (Grice1975-GRILAC-6, ), etc.), constitute the foundation of agents’ communication in MAS (berna2004communication, ; jang2005efficient, ; saunders1996evolution, ).
In MAS, illocutionary conversations are usually handled using descriptive and structured languages, such as JSON and RDF. Their purpose is to exchange information, not to perform or request actions. On the other hand, KQML (kqml1994, ), FIPA’s ACL, [^8] http://www.fipa.org/specs/fipa00061/XC00061D.html and rational programming and Agent-Oriented Programming (shoham1993agent, ) are examples of consolidated performative standards that describe agents’ beliefs, commitments, and actions. For instance, for an agent that exchanges a file with another and asks to summarise it, the MAS literature proposes using a structured language for the exchange and a performative query for the summarisation.
When it comes to LLMs, humans interact with them in free-text form; by extension, most MAS LLMs systems adopt natural language as the primary communication medium. While natural language captures the nuances of the principles mentioned above, it comes at the cost of complexity and ambiguity: an utterance such as "a car is coming" may range from being illocutionary (conveying information) to performative (conveying an implicit warning or an order). A few recent works in MAS LLMs propose to handle handshakes and errors with natural language communications; any other communication that routines and protocols can implement should otherwise favour structured languages (marro2024scalablecommunicationprotocolnetworks, ).
@@ 5 MAS LLMs do not Quantify the Emergence of Complex Behaviours
In many disciplines, such as system theory, economics, MAS, etc., an emergent behaviour is observed when a complex entity has properties or behaviours that its parts do not have on their own, and emerge only when they interact in a wider whole (altmann2024emergencemultiagentsystemssafety, ; fromm2005types, ). In the context of LLMs, emergence describes the increasing capabilities of a model at varying model size (brown2020languagemodelsfewshotlearners, ; kaplan2020scalinglawsneurallanguage, ). With MAS LLMs, behaviours that cannot be predicted via a static analysis of agents and their environment are also considered emergent Li_2023 ; park2023generative .
Emergent behaviours in MAS and MAS LLMs are often associated with open-ended environments, i.e., those MAS settings where LLMs interact and evolve freely. In an influential work park2023generative , Park et al. use LLMs to simulate a sandboxed society, with a focus on how single instructions influence the population and information spreads across the environment. While the authors show that simply nudging one agent causes other agents to engage in complex behaviours, the concept of emergence itself is never addressed formally.
@@
We thus surveyed papers published in the MAS LLMs and machine learning community between $2023$ and $2025$ that mention emergent behaviours to understand what methods and metrics employ to identify and quantify them. Out of more than $60$ papers analysed, only a few define clear metrics to measure emergent behaviours, while the majority qualitatively evaluate such behaviours and report the most notable. More details about the analysis and the list of papers are reported in Appendix C.
@@
In conclusion, in this body of research, we observe (i) no systematic definition of what constitutes an emergent behaviour, without reference to the system being analysed, therefore (ii) a lack of proper benchmarks to quantify them. In contrast, traditional MAS research generally insists on rigorously defined environments, quantifiable objectives, and formal verification methods. Emergence alone, especially when derived from loosely constrained prompts and undefined interaction dynamics, does not suffice to claim long-horizon MAS-level coordination or planning capabilities (turner2022formalizingproblemeffectregularization, ).
@@ 7 Conclusion
Acknowledgments
ELM was supported by the Alan Turing Institute. GLM was supported by UK Research and Innovation [grant number EP/S023356/1], in the UKRI Centre for Doctoral Training in Safe and Trusted Artificial Intelligence (www.safeandtrustedai.org). SM was supported by the EPSRC Centre for Doctoral Training in Autonomous Intelligent Machines and Systems n. EP/Y035070/1, in addition to Microsoft Ltd. MW was supported by an AI 2050 Senior Fellowship from the Schmidt Sciences Foundation. ELM, SM, and PT are affiliated with the Institute for Decentralized AI, which they thank for its support.